OpenStudy (anonymous):

integrate (4x^2-x+4)/(x^2+1)dx

OpenStudy (zarkon):

\[\frac{4x^2-x+4}{x^2+1}=4-\frac{x}{x^2+1}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is that the fraction simplified?

OpenStudy (zarkon):

yes...now use u substitution

OpenStudy (anonymous):

im not completely sure how you simplified it like that

OpenStudy (anonymous):

ah ok common denominators right

OpenStudy (zarkon):

\[\frac{4x^2-x+4}{x^2+1}=\frac{4x^2+4-x}{x^2+1}\] \[=\frac{4x^2+4}{x^2+1}-\frac{x}{x^2+1}\] \[=\frac{4(x^2+1)}{x^2+1}-\frac{x}{x^2+1}=4-\frac{x}{x^2+1}\]

OpenStudy (anonymous):

ok thx